# Robust Experimental Signatures of Phase Transitions in the Variational Quantum Eigensolver

### v2

@@ III.3 Noise Robust Detection of Phase Changes
Figure 4: The correlation functions $\braket{X_{0}X_{i}}$ and $\braket{Z_{0}Z_{i}}$ alongside energy and FS for $N=12$ at $B_{x}=0.1$ with open boundary conditions. The top row shows the ideal noise-free VQE correlation functions, while the middle row shows the experimental results on ibm_kyoto, starting from site $i=1$ on the $y$-axis. The $J_{2}/J_{1}$ axes are shared with Energy and FS plotted on the bottom rows. The FS is plotted at the midpoint between the two $J_{2}$ values used to calculate it. The differing number of peaks/transitions with respect to Figures 2 and 3 is due to these results being under open boundary conditions. All results were obtained using just TREX and Pauli twirling.
@@ IV Discussion and Outlook
While these results have shown that one can use NISQ devices to detect phases and their boundaries in VQE solutions, to be able to demonstrate a true advantage over classical computation requires finding VQE ground states for classically intractably large system sizes. Naturally this requires implementing a variational quantum algorithm optimization loop on real devices. Given the active research in compensating for problems of barren plateaus from overparameterization [63, 64] and noise [65, 66], our results suggest that characterization of complex zero temperature phase diagrams in large systems can be a real near term advantage for NISQ computation. Furthermore, while the interplay between EM and trainability of VQAs is an area of active research [67, 68, 16], fundamental limits on the ability of EM to fully compensate for noise of large system sizes have been derived [69, 70]. This raises the possibility that simple measures like the ones we propose, bolstered by stable EM methods like digitally shaping the noise via twirling and post-processing methods like TREX, could still retain use in the system size vs. error threshold range between EM and fully fault tolerant computation.
@@
In conclusion we found that by using a combination of simple target observables, we can reliably detect the properties of VQE ground state configurations in a noise robust manner on presently existing NISQ devices by interrogation of the experimental data from multiple angles, thus making the corruption of one observable by noise less detrimental. We demonstrated this utility experimentally by detecting rearrangement between phase domains for pre-optimized VQE solutions to a frustrated spin model. Our results have the potential for direct application to VQA solutions for medium and large sized systems, and could be useful even in cases where error mitigated observables are prohibitively expensive.
@@ Acknowledgments
Work at Los Alamos was supported by the LANL LDRD Program, the ASC Physics & Engineering Models Program, and in part by the Center for Integrated Nanotechnologies, a DOE BES user facility, in partnership with the LANL Institutional Computing Program for computational resources. We acknowledge the use of IBM Quantum services for this work. The views expressed are those of the authors, and do not reflect the official policy or position of IBM or the IBM Quantum team.

### v3

@@ III.3 Noise Robust Detection of Phase Changes
Figure 4: The correlation functions $\braket{X_{0}X_{i}}$ and $\braket{Z_{0}Z_{i}}$ alongside energy and FS for $N=12$ at $B_{x}=0.1$ with open boundary conditions. The top row shows the ideal noise-free VQE correlation functions, while the middle row shows the experimental results on ibm_kyoto, starting from site $i=1$ on the $y$-axis. The $J_{2}/J_{1}$ axes are shared with Energy and FS plotted on the bottom rows. The FS is plotted at the midpoint between the two $J_{2}$ values used to calculate it. The differing number of peaks/transitions with respect to Figures 2 and 3 is because the results plotted here are for open boundary conditions, while those are for periodic boundary conditions. All results were obtained using just TREX and Pauli twirling.
@@ IV Discussion and Outlook
While these results have shown that one can use NISQ devices to detect phases and their boundaries in VQE solutions, it is important to emphasize that the spirit of this work is not limited to the VQE algorithm, and can indeed be extended to any method which prepares ground state solutions to particular models. Going forward, to be able to demonstrate a true advantage over classical computation requires finding ground states for classically intractably large system sizes. In the case of the VQE, which is one of the most promising near term algorithms, this naturally requires implementing an optimization loop on real devices. Given the active research in compensating for problems of barren plateaus from overparameterization [76, 77] and noise [78, 79], our results suggest that characterization of complex zero temperature phase diagrams in large systems can be a real near term advantage for NISQ computation. Furthermore, while the interplay between EM and trainability of VQAs is an area of active research [80, 81, 21], fundamental limits on the ability of EM to fully compensate for noise of large system sizes have been derived [82, 83]. This raises the possibility that simple measures like the ones we propose, bolstered by stable EM methods like digitally shaping the noise via twirling and post-processing methods like TREX, could still retain use in the system size vs. error threshold range between EM and fully fault tolerant computation.
@@
In conclusion we found that by using a combination of simple target observables, we can reliably detect the properties of ground state configurations in a noise robust manner on presently existing NISQ devices by interrogation of the experimental data from multiple angles, thus making the corruption of one observable by noise less detrimental. We demonstrated this utility experimentally by detecting rearrangement between phase domains for pre-optimized VQE solutions to a frustrated spin model. Our results have the potential for direct application to circuits which prepare ground state solutions for medium and large sized systems, and could be useful even in cases where error mitigated observables are prohibitively expensive.
@@ Acknowledgments
Work at Los Alamos was supported by the LANL LDRD Program, and in part by the Center for Integrated Nanotechnologies, a DOE BES user facility, in partnership with the LANL Institutional Computing Program for computational resources. We acknowledge the use of IBM Quantum services for this work. The views expressed are those of the authors, and do not reflect the official policy or position of IBM or the IBM Quantum team.